# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

### v8

@@ 4.2. Results
Figure 5 compares each step of the text preprocessing pipeline. These steps are always applied sequentially in the pipeline. Beginning with $\varnothing$ (no text preprocessing), we see the performance of subsequent Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). For T and N, most of the data points above the equivalence line in precision, recall, and F1 score indicate that they can help syntactic OM. For R and S/L, some data points are above the equivalence line in recall, but most data points are below the equivalence line in precision and F1 score indicating that they do not play a positive role. In the violin plots: (a) Precision: The median increases with T and N but decreases with R and S/L. After T, the shape of the distribution is unchanged by R and S/L. (b) Recall: After T, the median increases slightly with each of N, R, and S/L. The shape of the distribution does not change after N. (c) F1 Score: The median increases with T and N but then decreases with R and S/L. The shape of the distribution does not change after N.
@@
Figure 5. Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). In the violin plots, three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3 (an artefact of the Seaborn (Waskom, 2021) package).
[…]
Comparison of text preprocessing pipeline.
  1. (1)
@@
    Figure 6 compares Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied. Most data points above the L = S line in precision and F1 score indicate that L is better than S in syntactic OM (assuming that post hoc correction is excluded). In the violin plots: (a) Precision: The median after L is greater than that achieved by S. The shape of the distribution is slightly different. (b) Recall: The median and the shape of the distribution are identical after S and after L. (c) F1 Score: The median after L is greater than for S. The shape of the distribution is identical.
    Figure 6. Stemming (S) vs Lemmatisation (L).
[…]
    Stemming (S) vs Lemmatisation (L).
  2. (2)
@@
    Figure 7 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL). For SP and SS, all data points located on the equivalence line indicate that there is no difference between SP and SS. For SP/SS and SL, most data points above the equivalence line in precision and F1 score indicate that SP/SS is more effective than SL. In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical in SP, SS, and SL.
@@
    Figure 7. Comparison of different stemmers: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
[…]
    Comparison of different stemmers.
  3. (3)
@@
    Figure 8 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT). All data points located on the equivalence line indicate that using POS Tagging in Lemmatisation does not help syntactic OM. In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
    Figure 8. Lemmatisation (L) without/with POS Tagging (LT).
[…]
    Lemmatisation (L) without/with POS Tagging (LT).
4.3. Implications
@@
For syntactic OM, the text preprocessing pipeline can be categorised into two phases. Phase 1 text preprocessing methods contain Normalisation and Tokenisation, whereas Phase 2 preprocessing methods consist of Stop Words Removal and Stemming/Lemmatisation. Phase 1 text preprocessing methods do not change word semantics, but change syntactic features such as formatting and typography. In contrast, Phase 2 text preprocessing methods change words to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
  1. (1)
@@
    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method. However, Phase 1 methods could also benefit from customisation in OM tasks. Some traditional methods originating from natural language processing (NLP) are not useful for OM and need to be adjusted. For example, sentence segmentation (i.e. splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations. Word tokenisation (i.e. breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of Pascal case and snake case.
  2. (2)
@@
    Phase 2 text preprocessing methods: We observe no benefit in using Phase 2 text preprocessing methods in syntactic OM. In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the matching performance. There are plausible explanations for this behaviour arising from the nature of ontologies as distinct from natural language text, as follows:
@@
      Stemming (S) vs Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs. On the other hand, S is more eager to find TPs, but eagerness can also lead to more FPs being found. Based on our experiments with no post hoc correction applied, L is better than S.
@@
      Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and we cannot see a significant difference between Porter and Snowball. Although the Lancaster Stemmer is more eager to detect more TPs, it does not lead to performance improvement as more FPs are also matched.
@@
  3. (3)
    In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods. This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement. There have been well-defined conventions established over the years, e.g. singularity, positive names, nouns for classes and verbs for properties (Schober et al., 2007; Taylor et al., 2015).
5. Text Preprocessing Pipeline Repair
The experimental results in Section 4 demonstrate that only Phase 1 text preprocessing (Tokenisation and Normalisation) helps with syntactic OM. The use of Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts. In this section, we explore the underlying cause of this behaviour and introduce our novel pipeline repair approaches.
5.1. Causal Analysis
  1. (1)
    Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning. For example, isReviewing is equivalent to is_reviewing. This means that applying Phase 1 methods only help to detect TPs, while the number of FPs remains unchanged. For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score.
    (1) $Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$
    (2) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
    (3) $F_{1}\ Score\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$
  2. (2)
    Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks. Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A). For example, isReviewing and isReviewedBy may be object properties with distinctly different meanings, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match. For this reason, recall is always increasing, but precision and the overall F1 score are less reliable.
    (4) $Prec\ ?=\frac{|A\cap R|}{|A|}=\frac{TP\uparrow}{TP\uparrow+FP\uparrow}$
    (5) $Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$
    (6) $F_{1}\ Score\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$
    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the condition to increase precision and F1 score is:
    (7) $\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$
@@ 5.1. Causal Analysis
5.2. Motivation
While Phase 2 text preprocessing methods have been shown to be less effective in syntactic OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs. As illustrated in Figure 9: (1) Phase 1 methods shift Alignment (A) towards Reference (R). The number of TPs increases, while the number of FPs decreases. (2) Phase 2 methods expand Alignment (A). The number of TPs increases, but the number of FPs also increases. (3) Our text preprocessing pipeline repair aims to collapse Alignment (A). The number of FPs decreases significantly, with a slight decrease in TPs.
Figure 9. Motivation of our context-based pipeline repair.
[…]
Motivation of our context-based pipeline repair.
5.3. Ad Hoc Logic-Based Repair
The ad hoc logic-based repair applies before the text preprocessing. One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing. The selection criteria are based on two ontology features: (1) entity names (class or property) are not duplicated within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (see Section 3). This approach is context-based because pairs of words may have the same or different meanings in different contexts. Based on these two features, we propose a simple Algorithm 1 to assemble the reserved word set. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively. For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. We define $S_{i}=\{i_{1},..,p,..,i_{n}\}$ is the string set of $e_{i}$ and $S_{j}=\{j_{1},...,p,...j_{m}\}$ is the string set of $e_{j}$. Let $p$ be the longest common substrings in both $S_{i}$ and $S_{j}$, then $S_{e}\div S_{j}=\{i_{1},..,i_{n},j_{1},...,j_{m}\}$. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we remove the words where $f(w)=w$ because their occurrence in the reserved word set has no effect on mapping results.
@@ 5.3. Ad Hoc Logic-Based Repair
  /* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$/
  /* Find duplicates in $O_{s}$/
@@
  /* Phase 2: Find immutable words in $S$/
@@
An example of generating and using a simple reserved word set is illustrated below. (1) The two object properties was_a_member_of and has_members from a single ontology have the same result “member” via traditional text preprocessing. Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities. The initial step (i.e. Phase 1 of Algorithm 1) is to add [“was”, “a”, “member”, “of”, “has”, “members”] to the reserved word set so that these two object properties would not have the same text preprocessing results. was_a_member_of preprocessed with skipping the reserved word set is “was a member of”, while has_members preprocessed with the reserved word set is “has members”. The revision step (i.e. Phase 2 of Algorithm 1) is to check whether there are immutable words in the reserved set. We can observe that the word “member” is immutable because it is the same before and after text preprocessing. Removing this word from the reserved word set still makes the two object properties different. Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”]. (2) The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies. For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical. While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping. As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
5.4. Post Hoc LLM-Based Repair
With strong background knowledge, LLMs can be helpful in detecting false mappings after text preprocessing. However, there is a wide range of LLMs with different prompting strategies that can be used in LLM-based repair. We aim to find a generic pipeline repair approach and the best practice to integrate LLMs in the repair phase. The experiment settings are as follows.
  1. (1)
    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED. We choose three datasets: WHOLE-FMA-NCI, WHOLE-FMA-SNOMED, and WHOLE-SNOMED-NCI.
  2. (2)

### v10

@@ IV-B Results
Fig. 6 compares each step of the text preprocessing pipeline. These steps are always applied sequentially in the pipeline. Beginning with $\varnothing$ (no text preprocessing), we see the performance of subsequent Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). For T and N, most of the data points above the equivalence line in precision, recall, and F1 score indicate that they can help OM. For R and S/L, some data points are above the equivalence line in recall, but most data points are below the equivalence line in precision and F1 score indicating that they do not play a positive role. In the violin plots: (a) Precision: The median increases with T and N but decreases with R and S/L. After T, the shape of the distribution is unchanged by R and S/L. (b) Recall: After T, the median increases slightly with each of N, R, and S/L. The shape of the distribution does not change after N. (c) F1 Score: The median increases with T and N but then decreases with R and S/L. The shape of the distribution remains unchanged after N, which also indicates that R and S/L do not help OM.
@@
Figure 6: Comparison of text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L). In the violin plots, three horizontal lines inside each violin plot show three quartiles: Q1, median, and Q3 (an artefact of the Seaborn [47] package).
@@
Fig. 7 compares Stemming (S) and Lemmatisation (L) after Tokenisation (T) and Normalisation (N) have been applied. Most data points above the L = S line in precision and F1 score indicate that L is better than S in OM (assuming that post hoc correction is excluded). In the violin plots: (a) Precision: The median after L is greater than that achieved by S. The shape of the distribution is slightly different. (b) Recall: The median and the shape of the distribution are identical after S and after L. (c) F1 Score: The median after L is greater than for S. The shape of the distribution is identical.
Figure 7: Comparison of Stemming (S) and Lemmatisation (L).
@@
Fig. 8 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL). For SP and SS, all data points located on the equivalence line indicate that there is no difference between SP and SS. For SP/SS and SL, most data points above the equivalence line in precision and F1 score indicate that SP/SS is more effective than SL. In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical in SP, SS, and SL.
@@
Figure 8: Comparison of different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL).
@@
Fig. 9 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT). All data points located on the equivalence line indicate that using POS Tagging in Lemmatisation does not help OM. In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
Figure 9: Comparison of Lemmatisation (L) without/with POS Tagging (LT).
IV-C Discussion
@@
For OM, the text preprocessing pipeline can be categorised into two phases. Phase 1 text preprocessing methods contain Normalisation and Tokenisation, whereas Phase 2 preprocessing methods consist of Stop Words Removal and Stemming/Lemmatisation. Phase 1 text preprocessing methods do not change word semantics, but change syntactic features such as formatting and typography. In contrast, Phase 2 text preprocessing methods change words to better capture semantic similarity, such as removing prefixes and suffixes or substituting common etymological roots.
@@
For Phase 1 text preprocessing methods, we observe that matching performance usually increases with each method. In most cases, OM only requires minor preprocessing using Phase 1 text preprocessing methods. This is because entity names in the ontology are often compound words that do not occur in natural language, but they are partially formalised by agreement. There have been well-defined conventions established over the years, e.g. singularity, positive names, nouns for classes and verbs for properties [48, 49]. However, Phase 1 text preprocessing methods could also benefit from customisation in OM tasks. Some traditional methods originating from natural language processing (NLP) are not applicable to OM and require adjustment. For example, sentence segmentation (i.e. splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations. Word tokenisation (i.e. breaking text into single words) could be rewritten to detect abbreviations commonly used in ontology concept names and to tackle the conflicting use of Pascal case and snake case.
@@
For Phase 2 text preprocessing methods, we observe no benefit in using these methods in OM. In some cases, Stop Words Removal and Stemming/Lemmatisation may even hamper the matching performance. There are plausible explanations for this behaviour arising from the nature of ontologies as distinct from natural language text, as follows:
@@
    Stemming (S) vs Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs. On the other hand, S is more eager to find TPs, but eagerness can also lead to more FPs being found. Based on our experiments, L is superior to S without post hoc correction.
@@
    Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and there is no significant difference between PS and SS. Although LS is more eager to detect more TPs, it does not lead to performance improvement as more FPs are also matched.
@@
V Text Preprocessing Pipeline Repair
The experimental results in Section IV demonstrate that only Phase 1 text preprocessing (Tokenisation and Normalisation) helps with OM. The use of Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation) does not improve performance and may even have negative impacts. In this section, we explore the underlying cause of this behaviour and propose our novel pipeline repair approaches.
V-A Causal Analysis
Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning. For example, isReviewing is equivalent to is_reviewing. This means that applying Phase 1 methods only help to detect TPs, while the number of FPs remains unchanged. For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score (Equations 12, and 3).
$Prec\uparrow=\frac{|A\cap R|}{|A|}=\frac{TP}{TP+FP}=1-\frac{FP}{TP\uparrow+FP}$ (1)
$Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (2)
$F_{1}\ Score\uparrow=\frac{2}{Prec\uparrow^{-1}+Rec\uparrow^{-1}}$ (3)
Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks. Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A). For example, isReviewing and isReviewedBy may be object properties with distinctly different meanings, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match. For this reason, recall is always increasing (Equation 5), but precision and the overall F1 score are less reliable (Equations 4 and 6).
[…]
$Rec\uparrow\ =\frac{|A\cap R|}{|R|}=\frac{TP\uparrow}{|R|}$ (5)
$F_{1}\ Score\ ?=\frac{2}{Prec\ ?^{-1}+Rec\uparrow^{-1}}$ (6)
If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then Equation 7 presents the condition to increase precision and F1 score.
$\frac{TP+\Delta TP}{TP+\Delta TP+FP+\Delta FP}>\frac{TP}{TP+FP}\Rightarrow\frac{\Delta TP}{\Delta FP}>\frac{TP}{FP}$ (7)
@@ V-A Causal Analysis
V-B Motivation
While Phase 2 text preprocessing methods have been shown to be less effective in OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs. As illustrated in Fig. 10: (1) Phase 1 methods shift Alignment (A) towards Reference (R). The number of TPs increases, while the number of FPs decreases. (2) Phase 2 methods expand Alignment (A). The number of TPs increases, but the number of FPs also increases. (3) Our text preprocessing pipeline repair aims to collapse Alignment (A). The number of FPs decreases significantly, with a slight decrease in TPs.
Figure 10: Motivation of our context-based pipeline repair.
V-C Ad Hoc Logic-Based Repair
The ad hoc logic-based repair is applied before text preprocessing. One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing. The selection criteria are based on two ontology features: (1) entity names (class or property) are not duplicated within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (see Section II). This approach is context-based because pairs of words may have the same or different meanings in different contexts. Based on these two features, we propose a simple Algorithm 1 to assemble the reserved word set. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively. For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. We define $S_{i}=\{i_{1},..,p,..,i_{n}\}$ is the string set of $e_{i}$ and $S_{j}=\{j_{1},...,p,...j_{m}\}$ is the string set of $e_{j}$. Let $p$ be the longest common substrings in both $S_{i}$ and $S_{j}$, then $S_{e}\div S_{j}=\{i_{1},..,i_{n},j_{1},...,j_{m}\}$. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we remove the words where $f(w)=w$ because their occurrence in the reserved word set has no effect on mapping results.
@@ V-C Ad Hoc Logic-Based Repair
/* Phase 1: Find duplicates in $O_{s}$ and $O_{t}$ */
 /* Find duplicates in $O_{s}$ */
@@
/* Phase 2: Find immutable words in $S$ */
@@
For example, consider two object properties was_a_member_of and has_members from a single ontology that have the same result “member” via traditional text preprocessing methods. Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities. The initial step (i.e. Phase 1 of Algorithm 1) is to add [“was”, “a”, “member”, “of”, “has”, “members”] to the reserved word set so that these two object properties would not have the same text preprocessing results. was_a_member_of preprocessed with skipping the reserved word set is “was a member of”, while has_members preprocessed with the reserved word set is “has members”. The revision step (i.e. Phase 2 of Algorithm 1) is to check whether there are immutable words in the reserved set. We can observe that the word “member” is immutable because it is the same before and after text preprocessing. Removing this word from the reserved word set still makes the two object properties different. Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”].
The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies. For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical. While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping. As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
We apply our context-based pipeline repair approach to the same OAEI datasets as above. Fig. 11 compares with and without context-based pipeline repair in 8 tracks with 49 distinct alignments. The text preprocessing pipeline methods implemented include less effective Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT). For precision, most data points are above the equivalence line, indicating that context-based pipeline repair significantly improves matching correctness. For recall, most data points are located on the equivalence line, and only a few data points are below the equivalence line, indicating that context-based pipeline repair slightly reduces matching completeness. Looking at F1 score, most of the data points are above the equivalence line, indicating that context-based pipeline repair also improves overall matching performance.
Fig. 11 also shows the relative improvement in matching performance made by Phase 2 preprocessing methods. This can be seen from the distance of data points from the equivalence line. We see that the matching performance improvement has Stemming (S) > Lemmatisation (L) > Stop Words Removal (R). Further, the matching performance improvement with different stemmers has Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS) and the matching performance improvement for lemmatisation (L) alone is very similar to Lemmatisation + POS Tagging (LT).
Figure 11: Context-based pipeline repair use in Phase 2 text preprocessing: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
Fig. 12 evaluates the scalability of ad hoc logic-based repair. The x-axis indicates the number of entities in each alignment, and the y-axis shows the number of reserved words per 100 entities. For all the alignments we tested, the value of reserved words per 100 entities is less than 10. With the number of entities increasing (i.e. the ontology size increases), the number of reserved words per 100 entities decreases. This may be because reserved words such as“has”, “is”, and “a” are commonly used in many ontologies and in multiple distinct entities within those ontologies. Conventionally, there are a limited number of such words, and therefore the reserved word feature of our approach is scalable in large-scale OM.
Figure 12: Reserved words per 100 entities compared to the number of entities.
V-D Post Hoc LLM-Based Repair
In this study, we aim to develop a generic pipeline repair approach and establish best practices for integrating LLMs in the repair phase by empirical analysis with selected OAEI tracks. LLMs vary with respect to prompting strategies that could be used in LLM-based repair.